Use the following information for questions 17 to 26 Egon Spengler is planning a charity benefit gala for his Save the CityCrash Duration Time Description Activ Cost/Week (weeks) Preceding Crashing Activity ity(ies) (weeks) Determine $900 gala themPurchase supplies, including Stay C Puff marshmallows $700 Pick musical act and sign contract 4 D $550 Sell tickets in $690 pWhat is the slack time for activity D? Enter your answer as a numerical whole number.What is the minimum completion time for the project after crashing (until no more crashing is possible)? 27 weeks 28 weeks 29If Egon has a budget of only $2,000 for crashing costs, what is the shortest possible completion time? 27 weeks 28 weeks 29 wHow much will it cost Egon to reduce the original total project time by three weeks? $1,530 $2,510 $3,550 $4,150How much did crashing that last (third) week cost Egon? $630 $900 $980 $1000
Answer #1

The critical path is the path which takes longest time to complete the project. Following is the calculation of critical path

A-B-D-G-J - 23

A-B-C-E-H-J - 30

A-B-C-E-I-J - 31 -----> Critical path

A-B-C-F-J - 27

A-B-C-E-I-J is a critical path because it takes maximum time to complete project.

a) Following is the calculation of slack

Task Predecessor Early Start DurationEarly Finish Late Finish Late Start Slaclk 12 19 18 25 26 12 19 19 26 26 12 12 12 19 19

Early Finish = Early Start + Duration

= 12 + 6

= 18

Late Finish = Early finish of Predecessor (Max time)

= 19

Slack = Late Finish - Early finish

= 19 - 18

= 1

The slack time of the activity D is 1

Following is the calculation of crashing project

Difference Crashing Time Optimal Duration Crashing (week) Activity (1) (2)Cost (4 Crashed Project Cost (Weeks) (week) (Week)

A-B-C-E-I-J is a critical path because it takes maximum time to complete project. Therefore, crashing of the project will be done for this path.

b) From the above table, the minimum completion time for the project after crashing is 31 - 3 = 28 weeks.

c) If the company has the budget of $2,000, then minimum completion of the project will be 31 - 2 = 29 weeks.

Difference Crashing Time Optimal Duration Crashing Activity (1) (2)Cost (4 Crashed Project Cost (Weeks) (week) (week) (Week)

d) The total cost for crashing the project is 900 + 980 + 630 = $2,510

e) Crashing the last week cost $630 (Activity E)